# Permutation twisted cohomology, remixed

### v2

@@ Introduction
    There is a homeomorphism between each $U(H)$ and the homogeneous spectrum of the $\mathbb{Z}$-graded endomorphism ring $\operatorname{End}_{\mathscr{L}(H)}^{\scriptscriptstyle\bullet}(k)$ in the localization $\mathscr{L}(H)$. In particular, $\operatorname{Spc}(\mathscr{K}(E))$ is a Dirac scheme.
In addition, under this open cover, there is an identification $\mathscr{K}(E)|_{U(1)}\cong\operatorname{D}_{b}(kE)$ - in other words, $U(1)$ is nothing more than the “cohomological open,” corresponding to the image under $\operatorname{Spc}$ of the localization $\mathscr{K}(E)\twoheadrightarrow\operatorname{D}_{b}(kE)$. We construct a refined open cover with these same properties.
Theorem B.
Let $G$ be a finite $p$-group. Then there exists an open cover
$\{U(H)\}_{H\in s_{p}(G)/G}$ (Construction 4.5) indexed by conjugacy classes of subgroups for which the following holds.
@@
    Each open $U(H)$ contains a unique closed point of $\operatorname{Spc}(\mathscr{K})$, $\mathfrak{m}_{H}$, corresponding to the unique closed point of $V_{{G}/\!\!/{H}}$; (Corollary 4.11)
@@
    The open $U(1)$ is equivalently the cohomological open $V_{G}=\operatorname{Spc}(\operatorname{D}_{b}(kG))$ of $\operatorname{Spc}(\mathscr{K}(G))$. In other words, we have an equivalence of categories $\mathscr{K}(G)|_{U(1)}\cong\operatorname{D}_{b}(kG)$; (Theorem 4.17)
@@
    Every endotrivial $C$ is a line bundle, i.e. over each open $U(H)$ is isomorphic to a shift of the tensor unit in the localization $\mathscr{K}(G)|_{U(H)}$. In particular, we have an isomorphism $C\cong k[h_{C}(H)]$ where $h_{C}$ denotes the unique h-mark function associated to $C$. (Theorem 4.13)
Finally, if $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, we completely recover the Balmer spectrum.
Theorem C. (Corollaries 6.4, 6.5).
If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, the comparison map is an open immersion and for each subgroup $H\leq G$, restricts to a local homeomorphism
$\operatorname{comp}_{\mathcal{L}_{G}(H)}:U(H)\to\operatorname{Spec}^{h}(\operatorname{End}_{\mathscr{L}_{G}(H)}^{\scriptscriptstyle\bullet}(k)).$
In this case $(\operatorname{Spc}(\mathscr{K}(G)),\mathcal{O}_{G}^{\scriptscriptstyle\bullet})$ is a Dirac scheme.
To obtain these results, it is seemingly necessary to twist by all effective endotrivials (or at minimum, a full-rank subset); see for instance Remark 4.12. In this sense, we believe our construction for $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is the “correct” generalization, and not too big. The not-so-secret sauce to the above results is a generalization of Balmer-Gallauer’s maps $a_{N}$ and $b_{N}$ for any effective endotrivial. For any effective endotrivial $C$, we construct unique (up to scaling) maps $\iota_{C}^{H}$ for every subgroup $H\leq G$ for which $\Psi^{H}(\iota_{C}^{H})$ is a quasi-isomorphism. Proving these maps exist is rather technical and relies heavily on the topological origins of effective endotrivials. These maps may not exist for chain complexes not arising from chain complexes of free modules over the orbit category $\Gamma_{G}$ (in particular, non-effective endotrivials), see Remark 3.8.
Theorem D. (Theorem 4.2).
Let $G$ be a finite $p$-group. For every effective endotrivial complex (see Definition 2.1) and subgroup $H\leq G$, there exists a homomorphism $\iota^{H}_{C}:k\to C[h_{C}(H)]$ such that $\Psi^{H}(\iota^{H}_{C})$ is a quasi-isomorphism.
Note that there are numerous maps that satisfy the above property (see Remark 4.3); our maps $\iota^{H}_{C}$ satisfy additional technical conditions. Finally, we prove some technical results regarding these chain complexes. Although they are not strictly necessary for our constructions, they contribute to our understanding of effective endotrivials. These propositions may be known to experts, but we reprove them for completeness.
Theorem E. (Proposition 3.11, Corollary 3.12).
Let $G$ be a finite group. Suppose $C$ is a bounded chain complex of permutation $kG$-modules arising from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules. Then $C$ has a contractible direct summand if and only if $C^{?}$ does. In particular, if $C$ is an indecomposable effective endotrivial, then $C$ arises from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules.
Open questions
There is but one essential ingredient missing to this story, that of noetherianity of the remixed twisted cohomology ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. At present, we conjecture that $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, and in fact, generated by homomorphisms $k\to C[s]$ for $C$ an irreducible endotrivial. Sadly, the strategy of [BG25, Lemma 12.12] to prove noetherianity seems to not carry over to the general setting. We (sketchily) remark Balmer-Gallauer’s proof can be adapted to any finite abelian $p$-group, but we omit a proof, as it follows near identically as in the elementary abelian case.
A new possibility we propose is comparing twisted cohomology to the usual cohomology ring. We construct a homomorphism from the twisted cohomology ring to a direct product of $p$-local group cohomology rings (see Definition 5.5), and whose kernel is nilpotent. Any group cohomology ring is well-known to be noetherian [Ven59, Eve61], therefore if the image of $\prod\hat{\Psi}$ is finite in each $p$-local group cohomology ring, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian modulo nilpotents. See Remark 6.6 for details.
Beyond this missing link, there remain numerous questions left to answer. First, it is quite reasonable to ask if a generalization of twisted cohomology exists for all finite groups. Not all endotrivials $C$ arise from a representation spheres (or any reasonable topological space), or can be expressed as (up to homotopy) a chain complex of permutation modules (see e.g. [BG23a, Corollary 5.6]). As a consequence, many of our maps $\iota^{H}_{C}$ cannot be constructed outside of the $p$-group setting; circumventing this issue is the main challenge. To go about constructing twisted cohomology, one may first have to determine $\operatorname{Pic}(\operatorname{K}_{b}(\operatorname{perm}(kG)))$, which at present is unknown (for non-$p$-groups). We ask if $\operatorname{Pic}(\operatorname{K}_{b}(\operatorname{perm}(kG)))$ and the subgroup of $\operatorname{Pic}(\mathscr{K}(G))$ generated by representation spheres are in fact equal, and if they are finite index subgroups of $\operatorname{Pic}(\mathscr{K}(G))$.
Second, while our twisted cohomology ring may be sufficient to capture the subtleties of $\operatorname{Spc}(\mathscr{K}(G))$, it is obviously more cumbersome. An explicit presentation in terms of generators and relations is at present seemingly beyond our reach. Even explicitly writing down endotrivials for larger groups (e.g. $p$-rank at least 3) is a significant challenge and relatively unexplored. Finally, although every endotrivial is a line bundle, it is currently unknown and an interesting question whether $\operatorname{Pic}(\mathscr{K}(G)|_{U(H)})\cong\mathbb{Z}$; we expect this to be the case. In correspondence, Gallauer gave a proof sketch of this for $p=2$ and $G$ elementary abelian, but not much is known beyond this case.
Organization
The paper is organized as follows. Section 1 covers some quick preliminaries regarding $p$-permutation modules, modular fixed points/Brauer quotients, and endotrivial complexes. Section 2 further reviews Borel-Smith functions. Section 3 covers chain complexes of free modules over the orbit category and effective endotrivials, and proves some technical lemmas. Section 4 constructs the local quasi-isomorphisms $\iota^{H}_{C}$ and the open cover $\{U(H)\}$ of the Balmer spectrum. Section 5 constructs the twisted cohomology ring. Section 6 proves injectivity of the comparison map and corollaries, generalizing the results of Balmer-Gallauer in the noetherian setting, and discusses noetherianity.
Notation
Our notation mostly follows [BG25], as opposed to [Mil25a]. Any group $G$ is assumed to be finite. The Weyl group G//H{G}/\!\!/{H} of $H\leq G$ is the subquotient $N_{G}(H)/H$. Given two subgroups $H,K\leq G$, we write $H=_{G}K$ if $H,K$ are $G$-conjugate, and $H\leq_{G}K$ to denote $H$ is a subgroup of a $G$-conjugate of $K$. We denote the set of all $p$-subgroups of $G$ by $s_{p}(G)$, and write $s_{p}(G)/G$ to denote a choice of conjugacy class representatives. We write $\operatorname{CF}(G)$ to denote the additive group of superclass functions, and if $G$ is not a $p$-group, we write $\operatorname{CF}(G,p)$ to denote the group of superclass functions valued on $p$-subgroups. We denote the trivial $kG$-module, the tensor unit of $\mathscr{K}(G)$, as $k$. Given a module $M$, $M[i]$ denotes the chain complex with $M$ in homological degree $i\in\mathbb{Z}$ and zero in all other degrees. [1]-[1] also denotes the shift functor on $\mathscr{K}(G)$. We denote by $\operatorname{H}^{\scriptscriptstyle\bullet}(G)=\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)$ the $\mathbb{N}$-graded cohomology ring of the tensor unit $k$ in $\operatorname{D}_{b}(kG):=\operatorname{D}_{b}(kG\operatorname{-mod})$ (with both odd and even-degree shifts for $p$ odd).
Acknowledgments
I am indebted to Paul Balmer and Martin Gallauer for numerous conversations and unwieldy email threads about their work and tensor-triangular geometry in abstracto. I also am grateful to Robert Boltje for his mentorship and suggestions, those of which heavily influenced Section 3 in particular, and the UC Santa Cruz Department of Mathematics for their generous graduate student travel support, which enabled conversations leading to the ideas in this paper to occur.
1. Preliminaries
We expediently review some of the essentials of permutation modules from [BG25] and endotrivial complexes [Mil25a]. For an overview of tensor-triangular geometry, we refer the reader to [Bal05, Bal10b], and for an in-depth “classical” overview of $p$-permutation modules and the Brauer quotient, we refer the reader to [Lin18, Chapter 5] or [Las23].

### v3

@@ Introduction
    $p$-groups of order up to $p^{3}$. (Corollary 7.9)
We briefly remark on the connection to [GAL25]. Gallauer considers twisted cohomology in an abstract context, and shows that when things behave particularly well, these constructions recover the Balmer spectrum and can be used to relate periodicity data between a tt-category and an associated twisted cohomology ring. The difficulty however, is showing such things behave well, and that is what we do here. In the language of [GAL25], finite generation of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ implies that the maps into the effective endotrivials, which Gallauer terms global sections, generate the topology of $\operatorname{Spc}(\mathscr{K})$. This determines a finite ample family of line bundles, which can be taken as the gradings which arise for a set of generators of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. This notion of ampleness almost recovers the notion of ampleness in algebraic geometry. If this holds, $\mathscr{K}(G)$ is a divisorial tt-category. We characterize the global sections for any effective endotrivial in 5.3.
Open questions
The most pertinent question presently is Noetherianity of the remixed twisted cohomology ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. We conjecture this holds.
Conjecture.
For any $p$-group $G$, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian. In particular, Theorem D holds in full generality for all $p$-groups.
Balmer–Gallauer showed this holds for $G$ elementary abelian [BG25, Lemma 12.12], and in Section 7 we adapt their techniques to show it holds in a wider berth of cases. However for certain groups of order $p^{3}$, we obtain the first example of a group for which twisted cohomology is not generated by its minimally-graded elements, suggesting that a list of generators of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ in general may grow rapidly as $|G|$ does. Our proofs rely very heavily on the topological origin of effective endotrivials, but unfortunately also rely on information about the classification of endotrivials specific to these groups. We propose a potential strategy, Remark 7.10, to prove Noetherianity in general, including a lemma which should be sufficient, but at present the full proof appears beyond our grasp. However, we have laid out the rest of the groundwork: if one can show that $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is Noetherian, then the analogues to the results of [BG25, Part II] hold for $G$.
While our twisted cohomology ring may be sufficient to capture the subtleties of $\operatorname{Spc}(\mathscr{K}(G))$, it is obviously cumbersome. An explicit presentation in terms of generators and relations, akin to the computations in [BG25, Section 17] are at present beyond our reach for arbitrary groups, but should be more feasible for the groups for which we’ve shown Noetherianity of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, as one now has an explicit list of generators.
Beyond twisted cohomology, there remain numerous questions. Perhaps the most immediate concerns the forerunner maps: our construction is ad-hoc and depend on a choice of direct sum decomposition, but we expect the existence of a canonical construction or description, perhaps arising from a topological point of view. Such a construction would resolve whether certain expected properties for forerunners hold, and elucidate the open covers we construct. We remark on this further, see Remark 4.4.
Our understanding of endotrivials, effective or not, is still incomplete, and this seems to be the limiting factor for proving Noetherianity. Though endotrivials are classified, explicitly writing them down for larger groups (e.g. non-abelian with $p$-rank at least 3) is a significant challenge. One fairly elementary question is as follows: in [MIL25b] it was shown that $\operatorname{Pic}(\mathscr{K}(G))$ is a biset functor in the sense of Bouc [BOU10], and “induction” of endotrivials is induced on the topological level by the norm map. How this presents on the level of $\mathscr{K}(G)$ seems difficult to determine, as it is (perhaps surprisingly) not given by tensor induction of chain complexes.
It is reasonable to ask if a generalization of twisted cohomology exists for all finite groups. Not all endotrivials $C$ arise from a representation spheres, so our forerunners cannot always be constructed outside of the $p$-group setting. Moreover, if $G/[G,G]$ is not a $p$-group, then not all endotrivials for $kG$ are line bundles. Explicitly, any shift of a $k$-dimension 1 $kG$-module $k_{\omega}$, with $\omega\in\operatorname{Hom}(G,k^{\times})$, is endotrivial. Related, we do not know a description of the subgroup of $\operatorname{Pic}(\mathscr{K}(G))$ generated by endotrivials induced from representation spheres; we expect this subgroup should be close to full-rank. We expect that one can construct an analogous open cover of $\operatorname{Spc}(\mathscr{K}(G))$ from this subgroup, and that some of the techniques of this paper carry over without much difficulty.
Finally, one may ask questions regarding descent with respect to the open cover $\{U(H)\}$ of $\operatorname{Spc}(\mathscr{K}(G))$ - in particular, there should be a descent spectral sequence associated to $\operatorname{Pic}(\mathscr{K}(G))$, which we will examine in forthcoming work. Although we prove every endotrivial is a line bundle, it is currently unknown and an interesting question whether $\operatorname{Pic}(\mathscr{K}(G)|_{U(H)})\cong\mathbb{Z}$ for all $H\leq G$; we expect this to be the case. In correspondence, Gallauer gave a proof sketch of this for $p=2$ and $G=V_{4}=C_{2}\times C_{2}$, but not much is known beyond this case.
Organization
The paper is organized as follows. Section 1 reviews tensor-triangular geometry, $p$-permutation modules and modular fixed points. Section 2 reviews endotrivial complexes and Borel-Smith functions. Section 3 covers chain complexes of free modules over the orbit category, effective endotrivials, and proves some technical results. Section 4 constructs the forerunners $\iota^{H}_{C}$ and the open cover $\{U(H)\}$ of the Balmer spectrum $\operatorname{Spc}(\mathscr{K}(G))$. Section 5 constructs the twisted cohomology ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ and proves compatabilities. Section 6 proves injectivity of the comparison map and corollaries. Finally, Section 7 proves Noetherianity of twisted cohomology for certain cases.
Notation
Our notation mostly follows [BG25], which we warn the reader differs considerably from [MIL24, MIL25a, MIL25b]. The Weyl group G//H{G}/\!\!/{H} of $H\leq G$ is the subquotient $N_{G}(H)/H$ of $G$. Given two subgroups $H,K\leq G$, we write $H=_{G}K$ if $H,K$ are $G$-conjugate, and $H\leq_{G}K$ to denote $H$ is a subgroup of a $G$-conjugate of $K$. We denote the set of all subgroups of $G$ by $s(G)$, the set of all $p$-subgroups of $G$ by $s_{p}(G)$, and write $s_{p}(G)/G$ to denote a choice of conjugacy class representatives. We write $\operatorname{CF}(G)$ to denote the additive group of superclass functions, i.e., functions $s(G)\to\mathbb{Z}$ constant on conjugacy classes, and if $G$ is not a $p$-group, we write $\operatorname{CF}(G,p)$ to denote the group of superclass functions valued on $p$-subgroups of $G$.
Throughout, $\mathscr{K}(G):=\operatorname{K}_{b}(p\operatorname{-perm}(kG))$. We denote the tensor unit of $\mathscr{K}(G)$ by $k:=k[0]$ for shorthand (as opposed to $\mathbb{1}$) and the tensor product by $\otimes$ (as opposed to $\otimes_{k}$). Given a module $M$, $M[i]\in\mathscr{K}(G)$ denotes the chain complex with $M$ in homological degree $i\in\mathbb{Z}$ and zero in all other degrees. We write $M:=M[0]$ if the context is clear. The shift functor on $\mathscr{K}(G)$ is also denoted [1]-[1]. We denote by $\operatorname{H}^{\scriptscriptstyle\bullet}(G)=\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)$ the $\mathbb{N}$-graded cohomology ring over $k$ in $\operatorname{D}_{b}(kG):=\operatorname{D}_{b}(\operatorname{mod}(kG))$ (including both odd and even-degree shifts for $p$ odd).
Acknowledgments
The author is indebted to Paul Balmer and Martin Gallauer for numerous conversations and unwieldy email threads about their work, comments on earlier versions of this work, and introducing the author to the world of tensor-triangular geometry. He also is grateful to Robert Boltje for his mentorship and suggestions during the course of the author’s Ph.D., and the UC Santa Cruz Department of Mathematics for their generous graduate student travel support, which enabled numerous conversations leading to the ideas in this paper to happen. Finally, he thanks Cat & Cloud Coffee of Santa Cruz, CA (on Swift St.) for their warm atmosphere, strong coffee, and impeccable breakfast burritos, countless of which were consumed during the lengthy inception of this project. During the course of this paper, the author was partially funded by a UC Santa Cruz dissertation year fellowship and an AMS-Simons travel grant.
1. Preliminaries: Tensor-triangular geometry and permutation modules
For the first two sections, we review the essential prerequisites of the paper. Section 1 is more on tensor-triangular geometric, and Section 2 concerns endotrivial complexes and their topological connections.
1.1. Tensor-triangular geometry
We begin by reviewing the basic concepts of tensor-triangular geometry. For more detailed overviews of tensor-triangular geometry, we refer the reader to [BAL05, BAL10b].
Definition 1.1.
Let $\mathscr{K}$ be an essentially small tensor-triangulated category. That is, $\mathscr{K}$ is an essentially small triangulated category with a symmetric monoidal structure such that for any $x\in\mathscr{K}$, the functor $x\otimes$- is an exact functor.
@@ 1.1. Tensor-triangular geometry
    A triangulated subcategory $\mathscr{I}\subseteq\mathscr{K}$ is thick if $\mathscr{I}$ is closed under direct summands, and is a thick $\otimes$-ideal if in addition, for all $x\in\mathscr{I}$ and $y\in\mathscr{K}$, $x\otimes y\in\mathscr{I}$.
@@
    A proper thick $\otimes$-ideal $\mathscr{P}\subset\mathscr{K}$ is prime if for all $x,y\in\mathscr{K}$, $x\otimes y\in\mathscr{P}$ implies $x\in\mathscr{P}$ or $y\in\mathscr{P}$.
@@
    Given a subset of objects $S\subset\mathscr{K}$, we let $\langle S\rangle$ denote the thick $\otimes$-ideal generated by $S$. This is equivalently the intersection of all thick $\otimes$-ideals containing $S$.
The following construction is due to Balmer [BAL05]. Let $\operatorname{Spc}(\mathscr{K})$ denote the set of prime thick $\otimes$-ideals of $\mathscr{K}$. For any $x\in\mathscr{K}$, let $\operatorname{supp}(x)\subseteq\operatorname{Spc}(\mathscr{K})$ denote the subset
$\operatorname{supp}(x):=\{\mathscr{P}\in\operatorname{Spc}(\mathscr{K})\mid x\not\in\mathscr{P}\},$
the support of $x\in\mathscr{K}$ Then the collection of all supports induces a closed basis on $\operatorname{Spc}(\mathscr{K})$, and a subset of $Z\subseteq\operatorname{Spc}(\mathscr{K})$ is of the form $Z=\operatorname{supp}(x)$ for some $x\in\mathscr{K}$ if and only if $Z$ is closed with quasi-compact complement [BAL05, Proposition 2.14]. We call $\operatorname{Spc}(\mathscr{K})$ the Balmer spectrum of $\mathscr{K}$. It is well-known that $\operatorname{Spc}(-)$ is contravariantly functorial [BAL05, Proposition 3.6] and the universal support datum for $\mathscr{K}$ in a precise sense [BAL05, Theorem 3.2]. In particular, any tt-functor (i.e., an exact $\otimes$-functor) induces a map on Balmer spectra.
Definition 1.2.
Recall that given any category $\mathscr{C}$, one can form its idempotent completion $\mathscr{C}^{\natural}$ (also called its Karoubi envelope), which roughly speaking, closes the category $\mathscr{C}$ under taking direct summands of objects. See for instance [BS01] for triangulated categories. Given an open $U\subseteq\operatorname{Spc}(\mathscr{K})$ with closed complement $Z=\operatorname{Spc}(\mathscr{K})\setminus U$, we set
$\mathscr{K}_{Z}:=\{x\in\mathscr{K}\mid\operatorname{supp}(x)\subseteq Z\},$
and define the localization
$\mathscr{K}|_{U}:=\left(\mathscr{K}/\mathscr{K}_{Z}\right)^{\natural}.$
Then $\mathscr{K}_{Z}$ is a thick $\otimes$-ideal of $\mathscr{K}$, $\mathscr{K}|_{U}$ is again a tensor-triangulated category, and we have a homeomorphism
$\operatorname{Spc}(\mathscr{K}|_{U})\cong U,$
see [BF07, Proposition 1.11]. This assignment endows $\operatorname{Spc}(\mathscr{K})$ with the structure of a locally ringed space [BAL05, Remark 6.4].
More generally, given any thick $\otimes$-ideal $\mathscr{I}\subseteq\mathscr{K}$ with canonical projection $\mathscr{K}\mapsto\mathscr{K}/\mathscr{I}$, the induced map on Balmer spectra $\operatorname{Spc}(\mathscr{K}/\mathscr{I})\to\operatorname{Spc}(\mathscr{K})$ induces a homeomorphism between $\operatorname{Spc}(\mathscr{K}/\mathscr{I})$ and the subspace $\{\mathscr{P}\in\operatorname{Spc}(\mathscr{K})\mid\mathscr{I}\subseteq\mathscr{P}\}$ of $\operatorname{Spc}(\mathscr{K})$ [BAL05, Proposition 3.11].
Example 1.3.
Let $G$ be a finite group (scheme), $k$ a field of characteristic $p$ prime, and $\operatorname{D}_{b}(kG)$ denote the bounded derived category of $kG$-modules. We have a homeomorphism
$\operatorname{Spc}(\operatorname{D}_{b}(kG))\cong\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)).$
This homeomorphism is realized by Balmer’s comparison map [BAL10a]
$\mathscr{P}\mapsto\langle\{\zeta\in\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)\mid\operatorname{cone}(\zeta)\not\in\mathscr{P}\}\rangle.$
We discuss an extended version of this comparison map in greater detail in the sequel. This theorem is due to Benson–Carlson–Rickard [BCR97] for finite groups (in very different language than we present) and was extended to finite group schemes by Friedlander–Pevtsova [FP07]. That the comparison map is a homeomorphism follows as a consequence of the classification of thick $\otimes$-ideals in both settings and Balmer’s classification theorem [BAL05, Theorem 4.10]. Surjectivity follows from Noetherianity of the cohomology ring [EVE61, FS97], and a general tt-geometric fact [BAL10a, Theorem 7.3], but as far as we are aware, there is no known direct proof showing the comparison map is even injective.
1.2. Permutation modules
Next, we review the basics of $p$-permutation modules. For a purely representation-theoretic overview of $p$-permutation modules and associate constructions, we refer the reader to [LIN18, Chapter 5] or [LAS23].